# QSTN: A Modular Framework for Robust Questionnaire Inference with Large Language Models

### v1

@@ Abstract
We introduce QSTN, an open-source Python framework for systematically generating responses from questionnaire-style prompts to support in-silico surveys and annotation tasks with large language models (LLMs). QSTN enables robust evaluation of questionnaire presentation, prompt perturbations, and response generation methods. Our extensive evaluation (>40>40 million survey responses) shows that question structure and response generation methods have a significant impact on the alignment of generated survey responses with human answers, and can be obtained for a fraction of the compute cost. In addition, we offer a no-code user interface that allows researchers to set up robust experiments with LLMs without coding knowledge. We hope that QSTN will support the reproducibility and reliability of LLM-based research in the future.
@@ 1 Introduction
To investigate some of these concerns, we present QSTN (pronounced ”Question”) - a Python framework designed to facilitate the execution of questionnaire-style experiments with LLMs. QSTN simplifies the process of creating robust variations of question prompts and answer generation methods, thereby facilitating reproducibility and the analysis of the reliability of LLM-based questionnaire research. QSTN provides a complete, modular pipeline for creating the questionnaire presentation, adjusting various parts of the prompt with perturbations, choosing the response generation method, performing inference, and finally, parsing the generated text.
@@ 2 Core Features
QSTN was developed with three objectives in mind: First, it enables robust evaluation of and with LLMs, addressing prompt sensitivity Tjuatja et al. (2024); Dominguez-Olmedo et al. (2024). QSTN is engineered to address this challenge directly through a highly modular and configurable design.
@@
Finally, QSTN is designed to be as easy to use as possible. Since we maintain the common prompt format of the system prompt and user prompt, adapting a project to QSTN is seamless. The package offers a complete pipeline from prompt creation and inference to parsing that can be done in only three function calls to the package. Integration with existing vLLM and OpenAI packages is straightforward.
@@ 2.1 Questionnaire Presentation
QSTN supports 3 questionnaire presentation modes:
@@ 2.2 Prompt Perturbation
Previous studies found that LLMs synthetic survey responses are highly sensitive to prompt perturbations and exhibit biases, such as token biases, recency bias, or A-bias Pezeshkpour and Hruschka (2024); Li and Gao (2025); Rupprecht et al. (2025); Dominguez-Olmedo et al. (2024); Röttger et al. (2024). QSTN can automatically randomize or reverse both the order of the questions within the survey and the order of answer options for each question to identify and mitigate these biases. This ensures that high performance is robust and independent of ordering. LLMs can be sensitive to small changes in prompt format He et al. (2024); Sclar et al. (2024). QSTN allows users to define custom answer label schemas (e.g., A/B/C, 1/2/3, i/ii/iii), enabling rigorous testing of a model’s robustness to superficial formatting changes. QSTN can perform the following Answer Option Perturbations:
  • Reversed Response Order: The order of answer options is reversed (e.g., a scale from ‘1: Very important‘ to ‘5: Not important‘ becomes ‘1: Not important‘ to ‘5: Very important‘).
@@
  • Odd/Even Scale Transformation: For scales with an even number of options, a semantically appropriate middle category is added, transforming it into an odd-numbered scale (e.g., by adding ‘Neutral‘). Conversely, for odd-numbered scales, we remove the middle category to create an even scale and adjust the integer label accordingly.
Figure 3: QSTN Supported Prompt Perturbations
@@ 2.3 Response Generation
While generative language models are designed to generate open-ended text, previous studies have implemented various approaches to constraining LLMs to closed-ended responses (e.g., Ma et al., 2024). We define Response Generation Methods as techniques used to elicit closed-ended responses from large language models to questionnaires Ahnert et al. (2025). QSTN supports the following Response Generation Methods:
@@
3 Using QSTN
The package containing QSTN can be installed in the desired environment using pip: pip install QSTN. QSTN is easily integrable into current workflows, needing just a total of three function calls for the most basic functionality, and it still allows users to freely define their prompts. A minimum usage example is given in Figure 5. By simply exchanging the function in the inference step, the questionnaire presentation can be adjusted, or a different type of parser can be selected. Additionally, building on this simple example, only one more module is needed to implement controlled prompt perturbations and response generation methods.
@@ Appendix A Questionnaire Presentation
Figure 6: Answer Distributions. Predictions and Ground Truth Distributions across the whole population compared for LLama-3.3-70B-Instruct and the question "How do you feel towards Gays and Lesbians?". We can see a clear shift towards the middle for this question, when models are given context of the previous questions and answers, which aligns more closely with human answers.
@@
We use the same prompt as that used in Bisbee et al. (2024), as displayed in Table 7. We adjust the output instructions to fit our choice response generation method and add all questions as instructions in the   battery presentation.

### v2

@@ Abstract
We introduce QSTN, an open-source Python framework for systematically generating responses from questionnaire-style prompts to support in-silico surveys and annotation tasks with large language models (LLMs). QSTN enables robust evaluation of questionnaire presentation, prompt perturbations, and response generation methods. Our extensive evaluation (>40 million survey responses) shows that question structure and response generation methods have a significant impact on the alignment of generated survey responses with human answers. We also find that answers can be obtained for a fraction of the compute cost, by changing the presentation method. In addition, we offer a no-code user interface that allows researchers to set up robust experiments with LLMs without coding knowledge. We hope that QSTN will support the reproducibility and reliability of LLM-based research in the future.
@@ 1 Introduction
To address and investigate some of these concerns, we present QSTN (pronounced “Question”) - a Python framework designed to facilitate the execution of questionnaire-style experiments with LLMs. QSTN simplifies the process of creating robust variations of question prompts and answer generation methods, thereby facilitating reproducibility and the analysis of the reliability of LLM-based questionnaire research. QSTN provides a complete, modular pipeline, as depicted in Figure 1, for creating the questionnaire presentation, adjusting various parts of the prompt with perturbations, choosing the response generation method, performing inference, and finally, parsing the generated text. We evaluate our framework on more than 40 million survey responses and find that the controlled variation of the experiment pipeline can increase the alignment of generated responses with human survey answers and that the responses can be obtained for a fraction of the compute cost.
@@ 2 Core Features
QSTN was developed with three objectives in mind: First, it enables robust evaluation of and with LLMs, addressing prompt sensitivity Tjuatja et al. (2024); Dominguez-Olmedo et al. (2024). QSTN is engineered to address this challenge directly through a highly modular and configurable design. Each part of the pipeline can be exchanged independently from the other parts.
@@
Finally, QSTN is designed to be as easy to use as possible. Since we maintain the common prompt format of the system prompt and user prompt, adapting a project to QSTN is seamless. The package offers a complete pipeline from prompt creation and inference to parsing, which can be done in only three function calls to the package. Integration with existing vLLM and OpenAI packages is straightforward.
@@ 2.1 Questionnaire Presentation
QSTN supports three distinct questionnaire presentation modes, as shown in Figure 2:
@@ 2.2 Prompt Perturbation
Previous studies found that LLMs synthetic survey responses are highly sensitive to prompt perturbations and exhibit biases, such as token biases, recency bias, or A-bias Pezeshkpour and Hruschka (2024); Li and Gao (2025); Rupprecht et al. (2025); Dominguez-Olmedo et al. (2024); Röttger et al. (2024). QSTN can automatically randomize or reverse both the order of the questions within the survey and the order of answer options for each question to identify and mitigate these biases. This ensures that high performance is robust and independent of ordering. Previous research has found that LLMs can be sensitive to small changes in prompt format He et al. (2024); Sclar et al. (2024). QSTN allows users to define custom answer label schemas (e.g., A/B/C, 1/2/3, i/ii/iii), enabling rigorous testing of a model’s robustness to superficial formatting changes. QSTN can perform the following Answer Option Perturbations, which are shown in Figure 3:
Figure 3: QSTN Supported Prompt Perturbations
  • Reversed Response Order: The order of answer options is reversed (e.g., a scale from ‘1: Very important’ to ‘5: Not important’ becomes ‘1: Not important’ to ‘5: Very important’).
@@
  • Odd/Even Scale Transformation: For scales with an even number of options, a semantically appropriate middle category is added, transforming it into an odd-numbered scale (e.g., by adding ‘Neutral’). Conversely, for odd-numbered scales, we remove the middle category to create an even scale and adjust the integer label accordingly.
@@ 2.3 Response Generation
While generative language models are designed to generate open-ended text, previous studies have implemented various approaches to constrain LLMs to closed-ended responses (e.g., Ma et al., 2024). We define Response Generation Methods as techniques used to elicit closed-ended responses from large language models to questionnaires Ahnert et al. (2025). QSTN supports the following Response Generation Methods, with examples being shown in Figure 4:
@@
3 Using QSTN
The package containing QSTN can be installed in the desired environment using pip. We support both a lightweight installation with pip install qstn, which only requires dependencies for API usage, and a full installation with pip install qstn[vllm] , which allows for local inference with vllm. QSTN is easily integrable into current workflows, requiring just a total of three function calls for the most basic functionality, and it still allows users to freely define their prompts. A minimum usage example is given in Listing 5. By simply exchanging the function in the inference step, the questionnaire presentation can be adjusted, or a different type of parser can be selected. Additionally, building on this simple example, only one more module is needed to implement controlled prompt perturbations and response generation methods.
@@ Appendix A Questionnaire Presentation
Figure 6: Answer Distributions. Predictions and Ground Truth Distributions across the whole population compared for LLama-3.3-70B-Instruct and the question “How do you feel towards Gays and Lesbians?”. We can see a clear shift towards the middle for this question, when models are given context of the previous questions and answers, which aligns more closely with human answers.
@@
We use the same prompt as that used in Bisbee et al. (2024), as displayed in Table 7. We adjust the output instructions to fit our choice response generation method and add all questions as instructions in the   battery presentation. For all models we use the default hyperparameter settings.